OpenStudy (anonymous):

Did I do this problem correctly??? log52x - 5 = -4 x=2.5

OpenStudy (anonymous):

log5(2x) - 5 = 4 Add 5 to both sides (log2x/log5) = 1 Cross multiply 5 up log2x = log 5 2x = 5 x = 5/2 5/2 = 2.5

OpenStudy (anonymous):

ok, thanks(: